# Adaptive Local Combining with Decentralized Decoding for Distributed Massive MIMO

### v1

@@ V Conclusion
$\begin{split}\!\!\!\!\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mt}\Big|^{2}\!\Bigg\}+\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{mt}\Big|^{2}\Bigg\}+2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mt}(\mathbf{v}_{ni_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{nt}\!\Bigg\},\end{split}$ (40)
The first term of (40) can be written as:
$\begin{split}\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!(\hat{\mathbf{g}}_{mt}+\tilde{\mathbf{g}}_{mt})\Big|^{2}\!\Bigg\}&=\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mt}\Big|^{2}\!\Bigg\}+\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mt}\Big|^{2}\!\Bigg\}\\ &\overset{(a)}{=}\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}\mathbb{E}\Big\{\!\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mt}\Big|^{2}\!\Big\}+\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mt}\!\Big\}\Big|^{2}\\ &-\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mt}\!\Big\}\Big|^{2}+\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}\mathbb{E}\Big\{\!\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mt}\Big|^{2}\!\Big\}\\ &+\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mt}\!\Big\}\Big|^{2}-\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mt}\!\Big\}\Big|^{2}\\ &\overset{(b)}{=}\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(A-L_{\mathcal{S}_{m}})\gamma_{mt}+\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Big|^{2}\\ &-\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Big|^{2}+\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(\beta_{mt}-\gamma_{mt})\\ &\overset{(c)}{=}\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Big|^{2}+\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(\beta_{mt}-\gamma_{mt})\end{split}$ (41)
we get (a) using same way as in (35) and (b) using the (10), (11) and (36). Similarly, the second term of (40) is first expanded same as in (a) in (41) using (35), then using the (12), (13) and (36), can be written as
$\begin{split}\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!(\hat{\mathbf{g}}_{mt}+\tilde{\mathbf{g}}_{mt})\Big|^{2}\!\Bigg\}&{=}\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(A+1)\gamma_{mt}+\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{A\gamma_{mt}}\Big|^{2}\\ &-\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\Big|a_{mt}\sqrt{A\gamma_{mt}}\Big|^{2}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(\beta_{mt}-\gamma_{mt})\\ &=\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{A\gamma_{mt}}\Big|^{2}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}\beta_{mt}.\end{split}$ (42)
The expectation in closed-form in third term of (40) can be written as:
$\begin{split}2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mt}(\mathbf{v}_{ni_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{nt}\!\Bigg\}=2\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Bigg)\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!a_{mt}\sqrt{A\gamma_{mt}}\Bigg).\end{split}$ (43)
Using (41), (42) and (43) in (40) and using this in (39), we get
$\mathbb{E}\Big\{|\text{BU}_{t}|^{2}\Big\}=|\text{DS}_{t}|^{2}+p^{u}_{t}\sum_{m=1}^{M}|a_{mt}|^{2}(\beta_{mt}-\delta_{mi_{t}}\gamma_{mt})-|\text{DS}_{t}|^{2}=p^{u}_{t}\sum_{m=1}^{M}|a_{mt}|^{2}(\beta_{mt}-\delta_{mi_{t}}\gamma_{mt}).$ (44)
The second term i.e., the pilot contamination term for the UE $t$ with the pilot sharing UE $k$ in (6), can be evaluated in closed-form as
$\begin{split}&\mathbb{E}\Big\{|\text{PC}_{tk}|^{2}\Big\}=p^{u}_{k}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mk}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{g}}_{mk}\Big|^{2}\Bigg\},\end{split}$ (45)
The RHS of (45) can be written as
$\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\Bigg\}+\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\Bigg\}+2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}(\mathbf{v}_{ni_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{nk}\!\Bigg\}$
Following the same steps as in (41), (42) and (43)
$\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!{\mathbf{g}}_{mk}\Big|^{2}\!\Bigg\}=\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}+\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(\beta_{mk}-\gamma_{mk}),$ (46)
$\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!{\mathbf{g}}_{mk}\Big|^{2}\!\Bigg\}{=}\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{A\gamma_{mk}}\Big|^{2}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}\beta_{mk},$ (47)
$2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}(\mathbf{v}_{ni_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{nk}\!\Bigg\}=2\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg)\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!a_{mt}\sqrt{A\gamma_{mk}}\Bigg),$ (48)
Therefore
@@
$\begin{split}&\mathbb{E}\Big\{|\text{PC}_{tk}|^{2}\Big\}=p^{u}_{k}\Big|\sum_{m=1}^{M}a_{mt}\sqrt{(A-\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}+p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{t}}\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (49)
The third interference term, for any non-pilot sharing UEs pair $t,k$, in (6) can be written as:
@@
$\begin{split}\mathbb{E}\Big\{|\text{UI}_{tk}|^{2}\Big\}&=p^{u}_{k}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mk}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{g}}_{mk}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\Bigg\}+p^{u}_{k}\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}+p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\tilde{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}\\ &+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\tilde{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}(1-\delta_{mi_{k}})\gamma_{mk}+p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}(\beta_{mk}-\gamma_{mk})\\ &+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}\gamma_{mk}+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}(\beta_{mk}-\gamma_{mk})\\ &=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{t}}\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (50)
The last term of interference in (6) can be written as:
$\begin{split}\mathbb{E}\Big\{|\text{GN}_{t}|^{2}\Big\}&=p^{u}_{k}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}+p^{u}_{k}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}\end{split}$ (51)
Using the (38), (44), (49), (50) and (51) in (6) gives the SINR closed-form expression in (14).
-B Proof of Lipschitz Continuity of $\nabla f(\bm{\delta}_{m})$:
@@
Consider the gradient component:
$g(\delta_{mi})=\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{mi}}=$ $\sum_{t=1}^{T}\frac{I_{mt}}{I_{mt}+S_{mt}}\frac{I_{mt}\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}$
$-\chi\left[2\lambda_{1}\max\left(0,\delta_{mi}-\delta^{2}_{mi}\right)\left(1-2\delta_{mi}\right)+2\lambda_{2}\max\left(0,\sum_{j=1}^{L_{p}}\delta_{mj}-A+1\right)\right]$
Boundedness Analysis:
Signal and Interference Terms: For $\delta_{mi}\in[0,1]$:
  • $S_{mt}>0$ and $I_{mt}>0$ (positive definite)
  • $S_{mt},I_{mt}\leq C_{1}$ (bounded above by network parameters)
  • $\frac{\partial S_{mt}}{\partial\delta_{mi}},\frac{\partial I_{mt}}{\partial\delta_{mi}}\leq C_{2}$ (bounded derivatives from (25) and (26))
Penalty Terms:

### v2

@@ V Conclusion and Future Work
$\begin{split}A_{1}&=\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!(\hat{\mathbf{g}}_{mk}+\tilde{\mathbf{g}}_{mk})\Big|^{2}\!\Bigg\}=\!\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\Big|^{2}\!\Bigg\}\!+\!\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\Big|^{2}\!\Bigg\}\\ &=\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}\!+\!\Big|\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\!\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}\\ &+\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|(a_{mt}\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}+\!\Big|\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2},\end{split}$
Using (10), (11) and (A2), we can rewrite $A_{1}$ as
@@
$\begin{split}A_{1}&=\Big|\!\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}+\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (A6)
Using (A1), the second term ($A_{2}$) of (A5) is expanded as:
$\begin{split}A_{2}&=\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\hat{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}\!+\!\Big|\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\!\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}\\ &+\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|(a_{mt}\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\tilde{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}+\!\Big|\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2},\end{split}$
Using the (12), (13) and (A2), we can rewrite $A_{2}$ as
$\begin{split}A_{2}&=\Big|\!\!\!\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{A\gamma_{mk}}\Big|^{2}\!\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}\!\!+\!\!\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!|a_{mt}|^{2}\beta_{mk}.\end{split}$ (A7)
Finally, the third term ($A_{3}$) of (A5) can be written as:
$\begin{split}A_{3}=2\Big(\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!a_{mt}\sqrt{(A{-}L_{\mathcal{S}_{m}})\gamma_{mt}}\!\Big)\Big(\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!a_{mt}\sqrt{A\gamma_{mt}}\Big).\end{split}$ (A8)
Using (A6), (A7) and (A8)
@@
$\begin{split}\mathbb{E}\Big\{\!|\text{UI}_{tk}|^{2}\!\Big\}&=\!\!\Big|\!\!\sum\limits_{m=1}^{M}\!\!\!a_{mt}\sqrt{(A{-}\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}\!\!\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\!\right|^{2}+\sum\limits_{m=1}^{M}\!|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{t}}\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (A9)
Using (A3), the term $\mathbb{E}\big\{|\text{GN}_{t}|^{2}\big\}$ of interference in (6) can be written as:
$\begin{split}&\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!|a_{mt}|^{2}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Big\}+\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!\!a_{mt}|^{2}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Big\}=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}\end{split}$ (A10)
Using the (A4), (A9) and (A10) in (6) gives the SINR closed-form expression in (14).
-B Proof of Lipschitz Continuity of $\nabla f(\bm{\delta}_{m})$
@@
The gradient component in (24) is $g(\delta_{mi})=\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{mi}}$ bounded as
  • $0<S_{mt},I_{mt}\leq C_{1}$ (always bounded)
  • $\frac{\partial S_{mt}}{\partial\delta_{mi}},\frac{\partial I_{mt}}{\partial\delta_{mi}}\leq C_{2}$ (bounded from (25) and (26))